Prepare a contribution margin income statement
x Ch 05 Ex 5-11 ents 6192-ACCT-2302., Help Save & Exit Submit Saved Ex 5-11 6 Check my work Exercise 5-11 Income reporting and break-even analysis LO P2 Blanchard Company variable costs are $176 per unit. The companys annual fixed costs are $664,400 manufactures a single product that sells for $220 per unit and whose total (1 Prepare a contribution margin income statement for Blanchard Company at the break-even point. BLANCHARD COM Contribution Margin Income Statement (at Break-Even) ences (2) Assume the companys fixed costs increase by $136,000. What amount of sales (in dollars) is needed to break even? Break Even Point in Dollars Choose DenominatorBreak Even Point in D Break-even point in dollars くPrey 10f 1lll Next>
Answer #1

Answer:

1

BLANCHARD COMPANY

Contribution Margin Income Statement (at Break-Even)

Amount

Percentage
of sales

Sales

3322000

100%

Variable cost

2657600

80%

Contribution margin

664400

20%

Fixed costs

664400

Net income

0

Working notes for the above answer is as under

Contribution margin ration

=(220-176)/220

=44/220

=20%

At break even point contribution margin is eqaqual to fixed cost so at break even point contribution margin is 664,400 equal to brake even point

So total sales

=664,400/20%

=332200

______________________________________________________________________

2

Break-even point in dollars

Choose Numerator:

/

Choose Denominator:

=

Break-even point in dollars

Total fixed costs

/

Contribution margin ratio

=

Break-even point in dollars

800400

/

20%

=

4002000